OpenStudy (anonymous):

write the slope intercept form of the equation of the line described. then find the distance between the point and the line. through(-5,4), perp. to y = 10/7x

OpenStudy (anonymous):

since it's perpendicular, then the slope of the line must be -7/10 (rule of perpendicular, it must be negative reciprocal) y = -7/10 x + b plug in your point 4 = (-7/10)(-5) + b 4 = 7/2 + b b = 1/2 y = (-7/10)x + 1/2